# PerlDiff: Controllable Street View Synthesis Using Perspective-Layout Diffusion Models

### v1

@@ Appendix A DDPM Preliminaries
Appendix B Implementation Details
Our PerlDiff utilize the pre-trained Stable Diffusion v1.4 [24], augmented with specific modifications to enhance scene control. Training was executed on a server equipped with eight Tesla V100 (32 GB) GPUs, across 60,000 iterations. An initial batch size of 16 was adjusted to a per-GPU batch of two for focused optimization, particularly for data samples comprising six view images per frame. The generation of samples conforms to the CFG rule [10], employing a guidance scale of 5.0 and the Denoising Diffusion Implicit Models (DDIM) [27] across 50 steps.
For scene manipulation, the text encoder within Stable Diffusion is retained, along with a weight-frozen CLIP to manage textual inputs and ConvNext for processing road maps. Feature extraction from Perl boxes is conducted via an MLP, optimized through Perl-based controlling module (Perl-CM) with randomly initialized weights. In contrast, certain modules inherit and freeze pre-trained weights from Stable Diffusion v1.4. The key parameters within Perl-CM, $\lambda_{b}$ and $\lambda_{s}$, are set to 5.0 to facilitate optimal image synthesis. Furthermore, DDIM [27] and CFG [10] are integrated into our training regimen, with a novel approach of omitting all conditions at a rate 10% to foster model versatility.
Optimization employs AdamW [17], without weight decay coefficient and a learning rate of $5\times 10^{-5}$, complemented by a warm-up strategy in the first 1,000 iterations. For additional insight into the realm of autonomous driving models, BEVFormer [14], StreamPETR [34], and CVT [41] were re-trained with original configurations tailored to our target resolution. The performance of BEVFusion [15] and MonoFlex [40] was evaluated using their provided codes and pre-trained weights.
Appendix C Additional Experiments
@@
  • Effectiveness of Controllable Generation on NuScenes (Subsection C.1)
  • Effectiveness of Classifier-Free Guidance Scale (Subsection C.2)
  • Effectiveness of View Cross-attention for Multi-View Consistency (Subsection C.3)
@@
C.1 Effectiveness of Controllable Generation on NuScenes
@@
Table 6: Controllability comparison for street view image generation on the NuScenes [1] validation set. A quantitative evaluation using 3D object detection metrics from BEVFusion [15].
@@
BEVControl* 27.14 13.19 19.91 0.94 0.34 0.96
PerlDiff (Ours) 26.61 24.69 30.71 0.82 0.28 0.76
C.2 Effectiveness of Classifier-Free Guidance Scale
@@
In Tab. 7, we assess the effect of the CFG [10] scale on the sampling of data generation. The term “scale” refers to the CFG scale, which is adjusted to balance conditional and unconditional generation. The transition from Method (b) to (e) indicates an increase in the CFG scale from 5.0 to 12.5. The results show an average increase of 2.60 in FID, an average decrease of 0.87% in mAP, an average reduction of 1.03% in NDS, a 0.02% increase in mAOE and a 1.07% drop in Vehicle mIoU. This provides substantial evidence that an excessively large CFG scale can degrade the quality of generated images and adversely affect various performance metrics.
@@
(a) 2.5 22.95 23.89 36.03 0.70 60.05 26.95
(b) 5.0 25.06 25.10 36.24 0.72 61.26 27.13
(c) 7.5 26.70 24.62 35.60 0.74 61.52 26.63
(d) 10.0 27.78 24.20 35.05 0.73 61.43 26.00
(e) 12.5 28.51 23.86 34.98 0.74 61.25 25.55
C.3 Effectiveness of View Cross-attention for Multi-View Consistency
View cross-attention ensures the seamless integration of visual data by maintaining continuity and consistency across the multiple camera feeds that are integral to current multi-functional perception systems in autonomous vehicles. Typically, autonomous vehicles feature a 360-degree horizontal surround view from a BEV perspective, resulting in overlapping fields of vision between adjacent cameras. Consequently, we facilitate direct interaction between the noise maps of each camera and those of the immediate left and right cameras. Given the noisy images from the current, left, and right cameras, designated as $\mathbf{Z}_{in}$, $\mathbf{Z}_{l}$, and $\mathbf{Z}_{r}$, respectively, the output of this multi-view generation is given by:
$\mathbf{Z}=\mathbf{Z}_{in}+\mathcal{C}(\mathbf{Z}_{in},\mathbf{Z}_{l},\mathbf{Z}_{l})+\mathcal{C}(\mathbf{Z}_{in},\mathbf{Z}_{r},\mathbf{Z}_{r}),$ (10)
where $\mathcal{C}(\cdot)$ represents the standard cross-attention operation, which accepts three input parameters: query, key, and value, respectively. This approach systematically integrates spatial information from various viewpoints, enabling the synthesis of images that exhibit visual consistency across distinct camera perspectives. Fig.7 offers a visual comparison of the model output with and without the application of view cross-attention.
Figure 7: Comparative visualization of outputs with (left) and without (right) view cross-attention. Red markers highlight discontinuities in the images generated without view cross-attention.
Appendix D Visualization Results
To further showcase the controllable generation capabilities of our PerlDiff, we present additional visual results. Fig. 8 provides extended examples of manipulating object orientation and size, while Fig. 9 demonstrates the precision of the model in positional control.
@@ Appendix D Visualization Results
Comparative visualizations, as illustrated in Fig. 10, reveal that while BEVControl* produces chaotic and indistinct attention maps leading to suboptimal image quality, our PerlDiff optimizes the response areas of the attention map, resulting in accurate object-level control information.
Additionally, Fig. 11 displays scene alterations by PerlDiff to mimic different weather conditions or times of day, showcasing its versatility in changing scene descriptions. Besides, Fig. 12 exhibits a series of synthetic samples generated with NuScenes validation set, showcasing the method’s adaptability and robustness in a variety of urban scenarios. Furthermore, Fig. 13 presents samples from KITTI [7] validation set, illustrating the application’s performance in real-world conditions.
Figure 8: Visualization of controlling object orientation and size. We present the results of our PerlDiff and BEVControl* in terms of controlling the orientation and size of objects. From top to bottom, the orientation of the object increases incrementally (-60 °, 60 °, and 90 °), and from left to right, we display objects of different sizes (car, truck, and bus).
Figure 9: Visualization of controlling object position. We present the results of PerlDiff and BEVControl* in terms of controlling the position of objects. From left to right, the objects progressively move forward, and from top to bottom, we display objects of varying sizes (car, truck, and bus).
@@
Figure 10: Visualization of cross-attention map results. From left to right, we present the generated images and corresponding cross-attention maps from our baseline BEVControl* and our PerlDiff.
@@
Figure 11: Qualitative visualizations on NuScenes [1]: day, night, and rain scenarios synthesized by our PerlDiff, exhibiting adaptability to various lighting and weather conditions.
Figure 12: Visualization of street view images generated by our PerlDiff on NuScenes [1] validation dataset. We show the ground truth ( $1^{st}$ row) and our PerlDiff ( $2^{nd}$ row).
@@
Figure 13: Visualization of street view images generated by our PerlDiff on KITTI [7] validation dataset. We show the ground truth (left) and our PerlDiff (right).

### v4

@@ A DDPM Preliminaries
Figure 5: Failure cases of PerLDiff, with red markers highlighting instances where, compared to the ground truth, PerLDiff generates images with the front and rear of vehicles reversed.
B Implementation Details
PerLDiff utilizes the pre-trained Stable Diffusion v1.4 [24], augmented with specific modifications to enhance scene control. Training was conducted on a server equipped with eight Tesla V100 (32 GB) GPUs over 60,000 iterations, which required two days. An initial batch size of 16 was adjusted to a per-GPU batch of two for focused optimization, particularly for data samples comprising six view images per frame. The generation of samples conforms to the CFG rule [9], employing a guidance scale of 5.0 and the DDIM [27] across 50 steps.
For scene manipulation, the text encoder within Stable Diffusion is retained, along with a weight-frozen CLIP to manage textual inputs and ConvNext for processing road maps. Feature extraction from PerL boxes is conducted via an MLP, optimized through PerL-based controlling module (PerL-CM) with randomly initialized weights. In contrast, certain modules inherit and freeze pre-trained weights from Stable Diffusion. The key parameters within PerL-CM, $\lambda_{b}$ and $\lambda_{s}$, are set to 5.0 to facilitate optimal image synthesis. Furthermore, DDIM [27] and CFG [9] are integrated into our training regimen, with a novel approach of omitting all conditions at a rate 10% to foster model versatility.
The optimization process employs AdamW [17] without a weight decay coefficient and with a learning rate of $5\times 10^{-5}$, complemented by a warm-up strategy during the first 1,000 iterations. BEVFormer [14], StreamPETR [34], and CVT [41] were retrained using original configurations tailored to our target resolution. The performance of BEVFusion [16] and MonoFlex [40] was assessed using their provided code and pre-trained weights.
C Limitation and Future Work.
Fig. 5 depicts several failure cases of PerLDiff, where the model erroneously generates vehicles with the front and rear orientations reversed, in contrast to the ground truth. This limitation arises from the usage of a PerL mask in PerLDiff, which does not account for the orientation on the 2D PerL plane. Future endeavors may explore video generation, extending to work such as DrivingDiffusion [12], Panacea [36], and Driving into the Future [35].
D Additional Experiments
@@
  • Effectiveness of Controllable Generation on NuScenes (Subsection D.1)
  • Effectiveness of Perl-based Cross Attention (Object) (Subsection D.2)
  • Effectiveness of View Cross-attention for Multi-View Consistency (Subsection D.3)
  • Effectiveness of PerLDiff Based on ControlNet (Subsection D.4)
  • Effectiveness of Classifier-Free Guidance Scale (Subsection D.5)
@@
D.1 Effectiveness of Controllable Generation on NuScenes
@@
Table 8: Controllability comparison for street view image generation on the NuScenes validation set. A quantitative evaluation using 3D object detection metrics from BEVFusion [16].
@@
BEVControl* 15.94 13.19 19.91 0.94 0.34 0.96
PerLDiff (Ours) 15.67 24.69 30.71 0.82 0.28 0.76
D.2 Effectiveness of Perl-based Cross Attention (Object)
To facilitate a better understanding of PerLDiff, we provide a detailed explanation of the PerL-based cross-attention (Object). As shown in Fig. 6, MagicDrive [6] utilizes text cross-attention from Stable Diffusion to implicitly learn a unified feature that concatenates text, camera parameters, and bounding boxes in the token dimension. In contrast, PerLDiff employs the PerL masking map as a prior, allowing each object condition to precisely control the corresponding pixel features. This results in more accurate positioning and orientation of objects in the generated images. Additionally, we integrated the object mask into the token dimension corresponding to the bounding box. As shown in Tab. 9, the results indicate improvements in BEVFormer, with NDS (e.g., 29.77 vs. 28.79 for MagicDrive) and mAOE (e.g., 0.73 vs. 0.81 for MagicDrive) demonstrating the effectiveness of PerLDiff in enhancing the performance of MagicDrive. Note that MagicDrive utilizes a single attention map for managing text, camera parameters, and boxes in the cross-attention process. Consequently, our ability to make improvements is constrained by the limited scope available for modifying the attention map within this architecture.
Figure 6: Overview of the PerL-based cross-attention (Object). MagicDrive employs text cross-attention to create a unified feature, while PerLDiff uses the PerL masking map to allow for precise control of pixel features for each object.
Table 9: Impact of integrating the PerL masking map (Object) into MagicDrive. We present the 3D object detection results based on BEVFormer [14], with outcomes showing superior performance emphasized in bold.
[…]
MagicDrive 15.92 15.21 28.79 0.81 0.57 0.95
MagicDrive + Mask 16.68 15.54 29.77 0.73 0.56 0.89
D.3 Effectiveness of View Cross-attention for Multi-View Consistency
View cross-attention ensures the seamless integration of visual data by maintaining continuity and consistency across the multiple camera feeds that are integral to current multi-functional perception systems in autonomous vehicles. Typically, autonomous vehicles feature a 360-degree horizontal surround view from a BEV perspective, resulting in overlapping fields of vision between adjacent cameras. Consequently, we facilitate direct interaction between the noise maps of each camera and those of the immediate left and right cameras. Given the noisy images from the current, left, and right cameras, designated as $\mathbf{Z}_{b}$, $\mathbf{Z}_{l}$, and $\mathbf{Z}_{r}$, respectively, the output of this multi-view generation is given by:
$\mathbf{\hat{Z}}=\mathbf{Z}_{b}+\mathcal{C}(\mathbf{Z}_{b},\mathbf{Z}_{l},\mathbf{Z}_{l})+\mathcal{C}(\mathbf{Z}_{b},\mathbf{Z}_{r},\mathbf{Z}_{r}),$ (11)
where $\mathcal{C}(\cdot)$ represents the standard cross-attention operation, which accepts three input parameters: query, key, and value, respectively. This approach systematically integrates spatial information from various viewpoints, enabling the synthesis of images that exhibit visual consistency across distinct camera perspectives. Fig.7 offers a visual comparison of the model output with and without the application of view cross-attention. Upon integrating view cross-attention into PerLDiff, the procedure of the PerL-CM is detailed in Algo. 1.
Figure 7: Comparative visualization of outputs with (Top) and without (Bottom) view cross-attention. Red markers highlight discontinuities in the images generated without view cross-attention.
Algorithm 1 PerL-based Controlling Module (PerL-CM)
2:Output: Updated $\mathbf{Z}$
5:$\mathbf{Z_{s}}\leftarrow\gamma_{s}\cdot\mathcal{A}_{s}\mathbf{H}_{m}+\mathbf{Z}$
6:$\mathcal{A}_{b}\leftarrow\textit{softmax}(\lambda_{b}\cdot\mathcal{M}_{b}+\mathbf{Z_{s}}\mathbf{H}_{b}^{T}/\sqrt{d})$
8:$\mathbf{Z_{b}}\leftarrow\gamma_{b}\cdot\mathcal{A}_{b}\mathbf{H}_{b}+\mathbf{Z_{s}}$
9:$\mathbf{\hat{Z}}\leftarrow\mathbf{Z}_{b}+\mathcal{C}(\mathbf{Z}_{b},\mathbf{Z}_{l},\mathbf{Z}_{l})+\mathcal{C}(\mathbf{Z}_{b},\mathbf{Z}_{r},\mathbf{Z}_{r})$
10:// maintain visual consistency via View cross-attention
11:$\mathbf{Z^{*}}\leftarrow\textit{softmax}(\mathbf{\hat{Z}}\mathbf{H}_{d}^{T}/\sqrt{d})\mathbf{H}_{d}+\mathbf{\hat{Z}}$
Table 10: Ablation study comparing PerLDiff with a ControlNet-based model. We present 3D object detection results based on BEVFormer, BEVFusion. Outcomes demonstrating superior performance are highlighted in bold.
[…]
ControlNet-based BEVFormer 20.46 18.07 28.48 0.87
GLIGEN-based 13.36 25.10 36.24 0.72
ControlNet-based BEVFusion 20.46 10.45 15.29 0.89
GLIGEN-based 13.36 15.24 24.05 0.78
Figure 8: Training curves of PerLDiff and ControlNet-based, illustrating that Ours converges more rapidly during training.
Figure 9: Qualitative comparison with MagicDrive. For scene controllability, PerLDiff demonstrates superior performance by results consistent with ground truth road information. Regions highlighted by yellow circles indicate areas where fail to align with ground truth.
D.4 Effectiveness of PerLDiff Based on ControlNet
In Tab. 10, we present an ablation study that replaces the architecture of PerLDiff with a ControlNet-based [39] model trained only on view cross-attention in Stable Diffusion. As shown in Tab.10, the performance of the ControlNet-based model is inferior to that of PerLDiff. Furthermore, Fig. 8 illustrates that PerLDiff employs a network architecture similar to GLIGEN [13], allowing it to converge more quickly on smaller datasets, such as NuScenes, compared to the ControlNet-based.
D.5 Effectiveness of Classifier-Free Guidance Scale
@@
In Tab. 11, we assess the effect of the CFG [9] scale on the sampling of data generation. The term “scale” refers to the CFG scale, which is adjusted to balance conditional and unconditional generation. The transition from Method (b) to (e) indicates an increase in the CFG scale from 5.0 to 12.5. The results show an average increase of 2.87 in FID, an average decrease of 0.87% in mAP, an average reduction of 1.03% in NDS, a 0.02% increase in mAOE and a 1.07% drop in Vehicle mIoU. This provides substantial evidence that an excessively large CFG scale can degrade the quality of generated images and adversely affect various performance metrics.
@@
(a) 2.5 12.36 23.89 36.03 0.70 60.05 26.95
(b) 5.0 13.36 25.10 36.24 0.72 61.26 27.13
(c) 7.5 15.52 24.62 35.60 0.74 61.52 26.63
(d) 10.0 16.32 24.20 35.05 0.73 61.43 26.00
(e) 12.5 16.86 23.86 34.98 0.74 61.25 25.55
Figure 10: Qualitative results of the generated images reveal discrepancies in background details. As indicated by the yellow circle, PerLDiff produces background elements that do not align with real images due to the incorporation of the PerL masking map.
E Visualization Results
@@
To further demonstrate the controllable generation capabilities of PerLDiff, we present additional visual results. Fig. 9 offers extended examples illustrating the superiority of PerLDiff in scene controllability, while Fig. 11 highlights its effectiveness in controlling object orientation. Fig. 12 reveals that BEVControl* produces chaotic and indistinct attention maps leading to suboptimal controllability, PerLDiff optimizes the response areas of the attention map, resulting in accurate object-level control.
Additionally, it is worth noting that, based on our experimental results, the key for temporal-based detection models lies in accurately positioning and categorizing objects in each frame; detailed information about objects, such as color and brand, is not crucial. As illustrated in Fig. 13, when provided with continuous frame inputs, the generated images by PerLDiff ensure that the positions and categories of objects, along with the road map, are consistently aligned with the specified conditions between adjacent frames.
Moreover, Fig. 14 displays scene alterations by PerLDiff to mimic different weather conditions or times of day, showcasing its versatility in changing scene descriptions. Furthermore, as illustrated in Fig. 10, PerLDiff generates background details that do not fully align with those of real images. This discrepancy arises because PerLDiff incorporates prior constraints to ensure accuracy in object detection, which can, in turn, negatively impact the fidelity of the background details.
Finally, Fig. 15 presents samples from KITTI validation set, illustrating the application’s performance in real-world conditions.
Figure 11: Qualitative comparison with MagicDrive. For object controllability, PerLDiff exhibits superior performance by generating objects at arbitrary angles. Regions highlighted by red circles denote scenarios where the images fail to achieve correct orientation.
@@
Figure 12: Visualization of cross-attention map results. From left to right, we present the generated images and corresponding cross-attention maps from our baseline BEVControl* and our PerLDiff. BEVControl* produces disorganized and vague attention maps, which result in inferior image quality. Conversely, PerLDiff method fine-tunes the response within the attention maps, resulting in more accurate control information at the object level and improved image quality.
Figure 13: Qualitative visualizations from the NuScenes. PerLDiff demonstrate consistent alignment of object positions and categories, along with the road map, when provided with continuous frame inputs, ensuring coherence between adjacent frames.
@@
Figure 14: Qualitative visualization on NuScenes demonstrating the effects of Text Cross-attention. From top to bottom: day, night, and rain scenarios synthesized by PerLDiff, highlighting its adaptability to different lighting and weather conditions.
@@
Figure 15: Visualization of street view images generated by our PerLDiff on KITTI validation dataset. We show the ground truth (left) and our PerLDiff (right).